# Quantum Algorithm of the GLMY Homology on Digraphs

### v1

@@ Abstract
Quantum algorithms for topological data analysis provide significant advantage over the best classical algorithm. Different from the previous simplical complex on points cloud, the GLMY homology introduced by Alexander Grigor’yan, Yong Lin, Yuri Muranov and Shing-Tung Yau [1], is defined on digraph and is a arising realm in Topological Data Analysis (TDA), which attracts more and more attention recently. We propose a quantum algorithm for the GLMY homology with significant advantage over the best classical algorithm. We design a universal encoding protocol for the quantum states and boundary operators of GLMY homology on digraphs. And a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm. The quantum algorithm for GLMY homology gives a quadratic speedup in general cases, and it gives an exponential quantum advantage in the case of the input data is given as a specification of paths.
@@ I Introduction
Topological Data Analysis (TDA) is a mathematical field focus on the ‘shapes’ of the data. By applying the topological tools, for example the simplicial homology theory and the knot theory etc., it is able to analyze complex high-dimensional data and identify key features of the data which may be difficult to be extracted with traditional methods. Quantum algorithm for TDA is first proposed by Lloyd, Garnerone, and Zanardi (LGZ) [2], and it gives significant advantage over the best classical algorithm. The foundations of the LGZ algorithm are linear-algebraic quantum machine learning (QML) algorithms [3, 4, 5, 6, 7, 8, 9, 10], which are one of the most rapidly growing fields in the research of quantum computing. Recent studies on quantum algorithms for TDA have achieved notable theoretical progress [11, 12, 13, 14, 15, 16, 17, 18]. The complexity of the LGZ algorithm is analyzed in [17] and the conclusion is that computing Betti numbers is NP-hard and it achieves quadratic speedup for inputs as vertices and edges. It can achieve exponential advantage only when inputs are given as a simplex list. The LGZ algorithms are also experimentally implemented on optical platform [19] and noisy quantum simulator [20].
@@
Here we propose a quantum algorithm for the GLMY homology on digraphs. The framework is based on the LGZ algorithm, which incorporates three components: the preparation of input states, Hamiltonian simulation of the Laplacian operator and phase estimate of the eigenvalues. We design a quantum state encoding protocol for digraphs based on paths in the GLMY homology. And we design a encoding protocol of the boundary map for quantum circuits. Our encoding protocol is universal for the paths on digraphs. Furthermore, a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm thus the norm matrix used in classical algorithm can be ignored. Given the input data as a specification of paths, the amplitude encoding method can be used to efficiently prepare the input state. The quantum algorithm for GLMY homology gives a quadratic speedup in general cases, and it gives an exponential quantum advantage in the special case of the input data is given as a specification of paths that can be sampled efficiently.
@@ III.1 Encoding of the Paths
In the case where we can efficiently sample $k$-paths, we propose that we can use the the amplitude encoding [34, 35, 36, 37] to prepare the input state $|P_{k}\rangle$. The amplitude encoding is previously used for data embedding in QML [38]. The amplitude encoding is to prepare the state
@@
where $N=\sum_{i}|c_{i}|^{2}$ is the normalization factor. In the case of preparing $|P_{k}\rangle$, the coefficients $c_{i}=1$ for all $i$, where $|i\rangle\in\Gamma_{k}$, and $\Gamma_{k}$ is the Fock space of molecular-system-like states with a conserved particle number. The scaling behavior for preparing $k$-path states is $n^{k}2^{k}/k!$, which is polynomially with $n$.
@@ III.3 Complexity Analysis
In the problem of solving the Betti number of simplicial complex is defined as: given a clique complex defined by its vertices and edges and an integer $k\geq 0$ as input, output the $k$th Betti number $\beta_{k}$ of the clique complex. This problem is proved to be NP-hard [17]. The mainly limiting factor for the complexity is the Grover search of the input state of the $k$-path, which is the same as the previous LGZ algorithm. The Grover search requires time complexity of $\Omega\big{(}\sqrt{\tbinom{n}{k+1}/\dim\Gamma_{k}|}\big{)}$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega(\sqrt{\dim\Gamma_{k}/\beta_{k}})$. The total runtime is lower bounded by $\Omega\big{(}\sqrt{\tbinom{n}{k+1}/\beta_{k}}\big{)}$.
However, in the case where we can efficiently sample $k$-paths, by using the amplitude encoding to prepare the input states, the quantum algorithm realizes an exponential speedup. The lower bound of the for preparing the state $\rho_{k}$ is $n^{k}2^{k}/k!$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega(\dim\Gamma_{k}/\beta_{k})$. Combining with these two lower bound, the total lower bound of the quantum algorithm for GLMY homology is $\Omega\big{(}n^{k}2^{k}\dim\Gamma_{k}/k!\beta_{k}\big{)}$, which is polynomial in $n$.
@@ IV Conclusion and Outlook
In this paper we propose the quantum algorithm for GLMY homology. We design a quantum state encoding protocol based on paths in the GLMY homology. And we design a efficient encoding protocol of the boundary map on quantum circuit. We prove that the null space of embedded homology equals to the null space of the the regular paths space projected on the embedded homology space. Finally, the quantum algorithm for GLMY homology gives an exponential quantum advantage in the case of the input states can be sampled efficiently, and it gives at least a quadratic speedup in other general cases.
@@ Acknowledgements.
This work is supported by Natural Science Foundation of Shandong Province (Grant No. ZR2025QC1476), National Key Research and Development Program of China (Grants No. 2021YFA1402104 and No.2021YFA0718302) and National Natural Science Foundation of China (Grant No. 12075310).

### v2

@@ Abstract
Quantum algorithms for topological data analysis provide significant advantage over the best classical algorithm. Different from the previous simplical complex on points cloud, the GLMY homology introduced by Alexander Grigor’yan, Yong Lin, Yuri Muranov and Shing-Tung Yau [1], is defined on digraph and is a arising realm in Topological Data Analysis (TDA), which attracts more and more attention recently. We propose a quantum algorithm for the GLMY homology with significant advantage over the best classical algorithm. We design a universal encoding protocol for the quantum states and boundary operators of GLMY homology on digraphs. And a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm. The quantum algorithm for GLMY homology provides a cubic speedup in general cases, and it can provide an exponential quantum advantage in the case of the input data is given as a specification of paths.
@@ I Introduction
Topological Data Analysis (TDA) is a mathematical field focus on the ‘shapes’ of the data. By applying the topological tools, for example the simplicial homology theory and the knot theory etc., it is able to analyze complex high-dimensional data and identify key features of the data which may be difficult to be extracted with traditional methods. Quantum algorithm for TDA is first proposed by Lloyd, Garnerone, and Zanardi (LGZ) [2], and it provides significant advantage over the best classical algorithm. The foundations of the LGZ algorithm are linear-algebraic quantum machine learning (QML) algorithms [3, 4, 5, 6, 7, 8, 9, 10], which are one of the most rapidly growing fields in the research of quantum computing. Recent studies on quantum algorithms for TDA have achieved notable theoretical progress [11, 12, 13, 14, 15, 16, 17, 18]. The complexity of the LGZ algorithm is analyzed in [17] and the conclusion is that computing Betti numbers is NP-hard and it achieves cubic speedup for inputs as vertices and edges. It can achieve exponential advantage only when inputs are given as a simplex list. The LGZ algorithms are also experimentally implemented on optical platform [19] and noisy quantum simulator [20].
@@
Here we propose a quantum algorithm for the GLMY homology on digraphs. The framework is based on the LGZ algorithm, which incorporates three components: the preparation of input states, Hamiltonian simulation of the Laplacian operator and phase estimate of the eigenvalues. We design a quantum state encoding protocol for digraphs based on paths in the GLMY homology. And we design a encoding protocol of the boundary map for quantum circuits. Our encoding protocol is universal for the paths on digraphs. Furthermore, a property of the GLMY homology is proved for the theoretical guarantee of the quantum algorithm thus the norm matrix used in classical algorithm can be ignored. In the case where the input data as a specification of paths is given as the input, the amplitude encoding method can be used to efficiently prepare the input state. The quantum algorithm for GLMY homology provides a cubic speedup in the case where the input is given as vertices and edges, and it can provide an exponential quantum advantage in the special case where the input data is given as a specification of paths.
@@ III.1 Encoding of the Paths
In the case where in the special case where the input data is given as a specification of paths, we propose that we can use the the amplitude encoding [34, 35, 36, 37, 38] to prepare the input state $|P_{k}\rangle$. The amplitude encoding is previously used for data embedding in QML [39]. The amplitude encoding is to prepare the state
@@
where $N=\sum_{i}|c_{i}|^{2}$ is the normalization factor. In the case of preparing $|P_{k}\rangle$, the coefficients $c_{i}=1$ for all $i$, where $|i\rangle\in\Gamma_{k}$, and $\Gamma_{k}$ is the Fock space of molecular-system-like states with a conserved particle number. The input state $|P_{k}\rangle$ can be prepared in $|\Gamma_{k}|^{2}(n\log n)^{2}$ steps [38], where $n\log n$ is the qubits number of the state $|P_{k}\rangle$.
@@ III.3 Complexity Analysis
In the problem of solving the Betti number of simplicial complex is defined as: given a clique complex defined by its vertices and edges and an integer $k\geq 0$ as input, output the $k$th Betti number $\beta_{k}$ of the clique complex. This problem is proved to be NP-hard [17]. The mainly limiting factor for the complexity is the Grover search of the input state of the $k$-path, which is the same as the previous LGZ algorithm. The Grover search requires time complexity of $\Omega\big(\sqrt{\tbinom{n}{k+1}/|\Gamma_{k}}|\big)$, and the lower bound of the phase estimation algorithm to compute the Betti number is $\Omega\big(\sqrt{|\Gamma_{k}|/\beta_{k}}\big)$. The total runtime is lower bounded by $\Omega\big(\sqrt{\tbinom{n}{k+1}/\beta_{k}}\big)$.
However, in the case where the input is given by the list of all $k$-path, by using the amplitude encoding to prepare the input states, the quantum algorithm can realize an exponential speedup. The complexity for preparing the state $\rho_{k}$ is $|\Gamma_{k}|^{2}(n\log n)^{2}$. In the case where $\Gamma_{k}$ is sparse, $|\Gamma_{k}|^{2}(n\log n)^{2}$ is scaling polynomially. Combining with the lower bound of the phase estimation, the total lower bound of the quantum algorithm for GLMY homology is $\Omega\big((n\log n)^{2}\sqrt{|\Gamma_{k}|^{5}/\beta_{k}}\big)$, which achieves an exponential speedup.
@@ IV Conclusion and Outlook
In this paper we propose the quantum algorithm for GLMY homology. We design a quantum state encoding protocol based on paths in the GLMY homology. We prove that the null space of embedded homology equals to the null space of the the regular paths space projected on the embedded homology space. Finally, the quantum algorithm for GLMY homology can provide an exponential quantum advantage in the case of the input states can be sampled efficiently, and it provides at least a cubic speedup in the case where the input is given as vertices and edges.
@@ Acknowledgements.
This work is supported by Science Challenge Project (Grant No. TZ2025017), National Key Research and Development Program of China (Grants No. 2021YFA1402104 and No.2021YFA0718302), National Natural Science Foundation of China (Grant No. 12075310) and Natural Science Foundation of Shandong Province (Grant No. ZR2025QC1476).